vidyo

New vidyo room:

Action Items:

News:

New indico page created.

Framework & production

Federica: last week we created image of a machine on cloud for Marco's group (big data tools available there now). We can read files from EOS public area, no need to copy the file locally.

Analysis

B physics

Alberto: Found a pb in the B MC. The Bs do not oscillate properly. Still under investigation. Franco: Luca staarted his thesis. He is now checking the MC samples. Alessio: P5' analysis: sent the comments to the referees.

Charged Higgs to top-beauty

Jacopo: submitted all the replies to the questions of the preapproval. got follow-up questions. Will reply asap. The MIT is checking the performances of a deep-neural-network wrt the BDT and is performing slightly worse.

Dibosons, VZ resonance (V=W,Z)

Jacopo: final reading of the 2l2q was last week. Went pretty well. Paul did an impressive job in answering all the Q from Bob Brown. Paul: Lisa's paper has been submitted to JHEP

4B full hadronic

Adres: Preapproves. ARC is reading, we expect some questions

Heavy New Quarks

Will update at the next meeting.

Higgs Dark Matter

Siew: moving away from this. I have a mangerial role now. Moving to the VH analysis we are about to start. We will add a link to the google doc of the analysis.

Same sign dileptons

Ugo: studying ttW and ttZ, in the new framework. Starting to look om 2017 dataset. Still missing the MC dataset for the same period.

Phase 2 physics studies

Pat: Duccio proposed an activity for the phase 2 tracker upgrade. HSCP is an analysis that will take advantage of the tracker upgrade. The new electronics will have only a threshold (and not 3 bits as before). Need to understand what is the impact of different thresholds on the HSCP analysis.

EPR

Pat: there are EPR for the upgrade. in the POGs and UPSG. 2.5 months EPR recognized for the muon ID studies. Siew: open to get EPR tasks.

Trigger(less)

Jacopo: Andrea started to reimplement the old electronics in P5. Will check the software w/ the firmware implemented by him. Around easter Andrea will be in Legnaro.

Detector

DT

Roberto: Useful meeting with the DT people.

Luminosity & Bril

Paul: CAlibrations and preparation for the beam. Had the ARC, waiting for a member, then we can go ahead with the approval.

Timing Layer

Tracker

A.O.B.:

Alessio: will have the reharsal of the thesis dissertation. -- Main.Franco Simonetto 04-10-2017 -- FrancoSimonetto - 2018-03-16
Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2018-03-19 - RobertoRossin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback